Consider the 2 ordered pairs (3,4) and (6,8).
(a) Find the slope of the line passing through these two points? I believe this is m=4/3
(b) Find the equation of the line passing through these two points. I believe this is y=4x/3
(c) Find the x and y intercepts of the line passing through these two points.
This is what I am stuck on. I can only find how to solve for the slope and y intercept and am getting confused.
Thank you.

Answer by Cromlix(4381)

Hi there,
a) You are correct m = 4/3
b) you are correct Equation of line = y = 4/3x
c) To find where a line cuts the x axis make y = 0
y = 4/3x
0 = 4/3x
x = 0
So (0,0)
To find where a line cuts the y axis make x = 0
y = 4/3x
y = 4/3(0)
y = 0
so (0,0)
line cuts at origin.
Hope this helps :-)